OpenStudy (anonymous):

1 + 3/5 + 5/8 + 3/8 + 36/113 + 3/11 divided by 6?

OpenStudy (phi):

1 + 3/5 + 5/8 + 3/8 + 36/113 + 3/11 I would do 1 + 5/8 + 3/8 first.= 1 + 8/8 = 1+1= 2 then 3/5 + 3/11 = 3*11/55 + 3*5/55 = 33/55 + 15/55 = 48/55 so you now have 2 + 48/55 + 36/113 add up, then multiply by 1/6 (which is the same as dividing by 6)

OpenStudy (anonymous):

3 509/1243 times 1/6 = 2119/3729 ?

OpenStudy (anonymous):

I did 2 + 48/55 + 36/113 first and then it would be 3 509/1243 times 1/6 = 2119/3729 right ?

OpenStudy (phi):

if we add 48/55 + 36/113 the common denominator is 55*113= 6215 multiply the first fraction by 113/113 and the second fraction by 55/55 you get 5424/6215 + 1980/6215 = (5424+1980)/6215 = 7404/6215 now add 2 which is (2*6215)/6215 = 12430/6215 we get 19834/6215 so 2 + 48/55 + 36/113 = 19834/6215 (this is definitely worth doing with a calculator to check the answer) now multiply by 1/6 \[ \frac{19834}{6215} \cdot \frac{1}{6} \] we can divide the top by 2 and the 6 by 2 to get \[ \frac{9917}{6215} \cdot \frac{1}{3} \\ \frac{9917}{18645}\]

OpenStudy (phi):

very ugly!